OpenStudy (anonymous):

Use matrix multiplication to find the reflection of v(-1,2) about the line y=x

OpenStudy (anonymous):

let the matrix be \[\left[\begin{matrix}0 & 1 \\ 1 & 0\end{matrix}\right]\] then...

OpenStudy (anonymous):

In general the reflection matrix about the line that forms angle alph with x axis is:\[A=\left[\begin{matrix}\cos2\alpha & \sin2\alpha\\ \sin2\alpha& -\cos2\alpha\end{matrix}\right]\] Hope this helps

OpenStudy (anonymous):

thanks for all but i know that ,idont how to use it ? if you can help me

OpenStudy (anonymous):

y=x forms an 45º angle with x axis. So cos90º = 0,sin90º=1 substitute this into A and you got it

OpenStudy (anonymous):

thanks myko

OpenStudy (anonymous):

\[reflection= \left[\begin{matrix}0 & 1 \\ 1 & 0\end{matrix}\right]\left(\begin{matrix}-1\\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

can you help me for (2,-5,3) please myko,and thanks for you

OpenStudy (anonymous):

(2,-5,3)?

OpenStudy (anonymous):

matrix multiplication to find the reflection of v=(2@-5@3) about the xz-plane

OpenStudy (anonymous):

General formula for reflection about ax+by+cz=0 plane is \[A=\left[\begin{matrix}1-2a ^{2} & -2ab & -2ac\\ -2ab & 1-2b ^{2}& -2bc \\-2ac & -2bc&1-2c ^{2}\end{matrix}\right]\] Can you find out by your self the equation of you plane and substitute values in the matrix?

OpenStudy (anonymous):

ok, equation of xz plane is y=0. So a=0, b=1,c=0. Just put this into the matrix and do like befor

OpenStudy (anonymous):

@m.f.h

OpenStudy (anonymous):

thanks for u myko